(12-5) exponent 0
Express in expanded form a exponent 5

Answer by Theo(13342)

if you mean (12 - 5) raised to the power of 0, the answer is equal to 1.
any expression raised to the power of 0 is equal to 1.
therefore (12 - 5) ^ 0 = 1
to expand this, you get (12 - 5) ^ 0 = 7 ^ 0 = 1
i think that's what you're looking for.
let me know if not.